The lateral edge of a pyramidal church spire is 61 ft. Each side of the

octagonal base is 22 ft. What will be the cost of painting the spire at

$0.025 a square foot?


1
Expert's answer
2021-11-29T11:50:29-0500

"\text{First, we calculate the total surface area of the octagon, where }\\
\text{Total surface area = Lateral surface area + base surface area}\\
\text{The base surface area is given by}\\
2(1+\sqrt{2})a^2=2(1+\sqrt{2})22^2\\
= 2336.96\\
\text{The perimeter of the base is given by $22\cdot8=176$}
\text{The lateral surface area is given by}\\
\frac12 \cdot 176 \cdot 61 = 5368\\
\text{Hence, the total surface area is }\\
5368 +2336.96= 7704.96sq. feet\\
\text{the cost of painting is }\\
7704.96 \cdot 0.025=\$192.624"